To standardize a solution of hydrochloric acid, you put 50.00 mL of it in a flask with a few drops of an indicator. You place 0.1524 M barium hydroxide in the buret. The buret reads 0.55 mL at the start and 33.87 mL when the equivalence point was reached. What is the molarity of the hydrochloric acid solution?
Answer #1

Concentration of Barium hydroxide, Ba(OH)2 = 0.1524 M = 0.1524 mol/L

Volume of Ba(OH)2 solution = (33.87 - 0.55) ml = 33.32 ml = 33.32 L / 1000 = 0.03332 L

Number of moles of Ba(OH)2 = 0.03332 L * 0.1524 mol/L = 0.005078 mol

Reaction :-

Ba(OH)2 (aq) + 2 HCl(aq) -----> BaCl2(aq) + 2 H2O (l)

From reaction, 1.0 mol of Ba(OH)2 required 2.0 moles of HCl so 0.005078 mol of Ba(OH)2 will produce 2.0 * 0.005078 mol = 0.010156 mol of HCl.

Volume of HCl required to reach the equivalence point = 50.00 ml = 50.00 L / 1000 = 0.05 L

Molarity of HCl solution = 0.010156 mol / 0.05 L = 0.2031 mol/L = 0.2031 M